OpenStudy (toxicsugar22):

8x^2+y^2-48+4y+68= write the equation in ellipse form

OpenStudy (amistre64):

8x^2 ... missing the other x value; is it the -48?

OpenStudy (amistre64):

8(x^2 -6x+___) +(y^2+4y+____) = -68 +____ + ____ 8(x^2 -6x+ 9) +(y^2+4y+4) = -68 +72 +4 8(x-3)^2 + (y+2)^2 = 8 8(x-3)^2 (y+2)^2 -------- + ------- = 1 8 8 (x-3)^2 (y+2)^2 -------- + ------- = 1 1 8

OpenStudy (toxicsugar22):

thanks

OpenStudy (toxicsugar22):

x^2+10y^2+4x+20y+4=0

OpenStudy (toxicsugar22):

write in ellipse form